OpenStudy (anonymous):

if i have 2y=-3x+6 how can i find slope and y-intercept? plzz help me

OpenStudy (anonymous):

Divide by 2 both side to get canonical form.

OpenStudy (anonymous):

First, get your equation into slope intercept form: y = -3/2 x + 3 (did this by dividing by 2)

OpenStudy (anonymous):

Slope : \[\text{slope}(f,x,x') = \frac{f(x) - f(x')}{x - x'} \]

OpenStudy (anonymous):

But here you just have to take y = slope x + offset.

OpenStudy (anonymous):

Or in non math language, slope = rise/run

OpenStudy (shadowfiend):

Yes, once you've divided by 2 on both sides you'll have y = mx + b where m is your slope and b is your y-intercept. Magic! :)